OpenStudy (anonymous):

can anyone help me with this! The height h in feet of a baseball on Earth after t seconds can be modeled by the function h(t) = -16(t – 1.5)2 + 36, where -16 is a constant in ft/s2 due to Earth's gravity. The gravity on Mars is only 0.38 times that on Earth. If the same baseball were thrown on Mars, it would reach a maximum height 59 feet higher and 2.5 seconds later than on Earth. Write a height function for the baseball thrown on Mars.

OpenStudy (anonymous):

@Kstate

OpenStudy (anonymous):

@Gixxer_J

OpenStudy (anonymous):

I cant help you sorry

OpenStudy (gixxer_j):

i feel stupid i dont know this :(

OpenStudy (anonymous):

Thanks for trying hun.

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@ahvei can you help her please

OpenStudy (anonymous):

they are multiple choice A. Y= -6.08(t-4)^2+95 B. Y=-6.08 (t-2.5)^2+59 C. Y= -0.38 (t-2.5)^2+59 D. Y= -0.38 (t-4)^2+95

OpenStudy (anonymous):

@Kstate

OpenStudy (anonymous):

I think it is A but im not sure